OpenStudy (anonymous):

Kendra wants to be able to make withdrawals of $60,000 a year for 30 years after retiring in 35 years. How much will she have to save each year up until retirement if her account earns 7% interest?

OpenStudy (anonymous):

I am just wondering is this an inequality?

OpenStudy (anonymous):

Ok so this is what I believe it is. We need to make this into an equation I believe this is a greater than or equal to problem. So I made an equation 35(1.07x)>or=60000*30 1.07 for interest 37.45x=1800000 simplified You want x alone so you divide 37.45 from each side 37.45x=1800000 /37.45 /37.45 x=48064.09 So she would need to save $48,064.09

OpenStudy (anonymous):

Do you understand this?